# Jaime Angulo Pava  and  Alexander Muñoz

### v3

@@ Abstract
The aim of this work is to study the Airy and Schrödinger operators on looping-edge graphs, a graph consisting of a circle and a finite number $N$ of infinite half-lines attached to a common vertex. Characterizations of extensions generating unitary and contractile dynamics are provided for the Airy operator in terms of self-orthogonal subspaces and linear operators acting on indefinite inner product spaces (Krein spaces) associated to the boundary values at the vertex. Employing the same abstract techniques used in the former case, we characterize the self-adjoint extensions of the Schrödinger operator on $\mathcal{T}$-shaped graphs, i.e., to graphs that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$. The theory established in this manuscript can be considered as a first step towards a better understanding of the dynamics for the remarkable dispersive models NLS and KdV on metric graphs.
@@ 1. Introduction
Nonlinear evolution models on metric graphs arise as models in wave propagation, for instance, in quasi-one-dimensional systems (e.g., meso- or nanoscale) that resemble a thin neighborhood of a graph. These models provide a convenient means to study various physical effects in the real world, both from theoretical and experimental perspectives. The flexibility in setting the geometry of the graph configuration allows for the creation of different dynamics. To mention a few examples of bifurcated systems: Josephson junction structures, networks of planar optical waveguides and fibers, branched structures associated with DNA, blood pressure waves in large arteries, nerve impulses in complex arrays of neurons, conducting polymers, and Bose-Einstein condensation (see [17, 19, 20, 24, 25, 27, 30, 35, 37, 40, 45] and reference therein).
Recently, nonlinear models on graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries model, have been studied extensively (see [2, 11, 12, 8, 9, 13, 14, 15, 3, 30, 40] and references therein). From a mathematical viewpoint, an evolution model on a graph is equivalent to a system of PDEs defined on the edges (intervals), where the coupling is determined exclusively by the boundary conditions at the vertices (known as the ”topology of the graph”), which governs the dynamics on the network via groups. This area of study has attracted significant attention, particularly in the context of soliton transport. Solitons and other nonlinear waves in branched systems provide valuable insights into the dynamics of these models.
@@
We recall that a metric graph $\mathcal{G}$ is a structure represented by a finite number of vertices $V=\{\nu_{i}\}$ and a set of adjacent edges at the vertices $E=\{e_{j}\}$ (for further details, see [17]). Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$. Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical ”wires” or ”networks” connecting them. The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$. This identification introduces the coordinate $x_{e}$ along the edge $e$.
An example of a metric graph, which will be the focus of our study here, is a looping-edge graph. This is a graph composed of a circle and a finite number $N$ of infinite half-lines attached to a common vertex $\nu=L$. If the circle is identified with the interval $[-L,L]$ and the half-lines with $[L,\infty)$, we obtain a particular metric graph structure that we denote with $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$ and $V=\{L\}$ (see Figure 1). We use the subscripts 00 and $j$ to refer to the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$, $j=1,\dots,N$, respectively. The special case $N=1$ is known as a tadpole graph (see also Figure 1).
$-L$
@@
A wave function U defined on $\mathcal{G}$ will be understood as an $(N+1)$-tuple of functions $\textbf{U}=(\phi,(\psi_{j})_{j=1}^{N})=(\phi,\psi_{1},\dots,\psi_{N})$, where $\phi$ is defined on $e_{0}=[-L,L]$ and $\psi_{j}$ is defined on $e_{j}=[L,\infty)$ for $j=1,\dots,N$.
@@
$i\textbf{U}_{t}\ +\Delta\textbf{U}+(p+1)|\textbf{U}|^{2p}\textbf{U}=\textbf{0},\quad p>0,$ (1.1)
@@
$-\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
@@
Since the deficiency indices of $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ are equal to $2+N$, the extension theory of Krein and von Neumann provides a $(2+N)^{2}$-parameter family of self-adjoint extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ of $\mathcal{H}_{0}$ on $L^{2}(\mathcal{G})$, where the action of $-\Delta_{ext}$ is given by $-\Delta_{ext}=-\Delta$. Each member of this family defines a unitary dynamic for the linear evolution problem
$\left\{\begin{array}[]{ll}i\textbf{U}_{t}=-\Delta_{ext}\textbf{U}\\ \textbf{U}(0)=\textbf{u}_{0}\in D(-\Delta_{ext}).\end{array}\right.$ (1.4)
We point out that our approach, based on Krein spaces [44], demonstrates that every extension operator $(-\Delta_{ext},D(-\Delta_{ext}))$ can be characterized by the boundary conditions at the vertices $\nu=\pm L$, treating the point $-L$ as a terminal vertex through a relatively simple matrix process (see Theorems 5.2-5.3 below). At this level, our approach can be seen as a complement to the works of Exner et al. [27, 28, 29], which, to our knowledge, has not formally appeared in the literature on $\mathcal{G}$. Furthermore, in general, the extensions $(-\Delta_{ext},-D(\Delta_{ext}))$ do not necessarily satisfy continuity conditions at the boundary of $[-L,L]$. (see, for instance, Example 8 associated to $\delta^{\prime}$- type interactions on looping-edge graphs and Example 10 below by considering $-L$ as a terminal vertex). Moreover, our approach to obtaining these extensions will be based on those associated with the Airy operator (1.7) which is the main focus of our study. Both extensions theories to be established can be seen as a first step for future studies on the existence and stability of soliton profiles for both NLS-(1.1) and KdV-(1.8) models, for instance, on a looping-edge graph or on a $\mathcal{T}$-shaped graph, i.e., a graph that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$ (see Figure 4). In the case of looping-edge graphs, the extensions must ensure that the wave functions are continuous at least on the loop (i.e., $\phi(L)=\phi(-L)$). For example, in the study of electron motion in thin metallic ”wires,” it is required that the wave function remains continuous at the junction $\nu=L$. These boundary conditions have a clear physical interpretation, representing the conservation of probability flow at the junction (see [28]).
@@
$\begin{array}[]{ll}D_{Z,N}=\{\mathbf{U}\in H^{2}(\mathcal{G}):&\phi(L)=\phi(-L)=\psi_{1}(L)=\cdot\cdot\cdot=\psi_{N}(L)\;\text{and}\\ &\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{j=1}^{N}\psi_{j}^{\prime}(L+)+Z\psi_{1}(L)\},\end{array}$ (1.5)
where for any $n\geqq 0$,
@@
The boundary conditions in (1.5) are called $\delta$-type if $Z\neq 0$, and Neumann-Kirchhoff type if $Z=0$. The parameter $Z$ is a coupling constant between the loop and the several half-lines (see Example 7 and Remark 1). The choice of coupling at the vertex $\nu=L$ corresponds to a feasible quantum-wire experiment (see [27], [29], and references therein). It should be noted that, based on the domain $D_{Z,N}$, several recent studies have been conducted, particularly on the existence and stability of standing-wave solutions for the model in (1.1) (see [3, 4, 5, 6, 7, 16, 21, 22, 34, 42, 41], and references therein).
Our second focus in this manuscript is associated to obtain a characterization of all boundary conditions under which the Airy-type evolution equation
$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t),\quad x\neq 0,t\in\mathbb{R},\textbf{e}\in E,$ (1.6)
with real coefficients given by the sequences $(\alpha_{\textbf{e}})_{e\in E}$, $(\beta_{\textbf{e}})_{e\in E}$, generates either a contraction semigroup or a unitary group on a looping edge graph (see Theorems 4.2, 4.3 and 4.7). Our approach is based on the abstract results in [44] applied to the case of the Airy-operator
@@
when we consider the metric graph $\mathcal{G}$ represented by $E$ defined above and treating the point $-L$ as a terminal vertex (as in the case of a $\mathcal{T}$-graph). Therefore, notions of self-orthogonal subspaces and operators based on indefinite inner product spaces (Krein spaces) will be necessary. Thus, we get in particular a description of all (skew-)self-adjoint realizations of $A_{0}$, - i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$. For instance, in Example 1 below we get $\delta$-interactions type boundary condictions on a looping-edge graph in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [8]) under conditions on $\alpha_{\textbf{e}}$ and $\beta_{\textbf{e}}$. Also, we establish conditions for obtaining extensions of $A_{0}$ generating contractile dynamics.
@@
We recall that the evolution linear model in (1.6) is associated to the following vectorial Korteweg-de Vries (KdV) equation
$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t)+2u_{\textbf{e}}(x,t)\partial_{x}u_{\textbf{e}}(x,t),$ (1.8)
@@
on the metric graph $\mathcal{G}$. Recently, Mugnolo, Noja, and Seifert [38] characterized all boundary conditions under which the Airy equation in (1.6) generates either a semigroup or a group on a metric graph that has a structure represented by the set $E\equiv E_{-}\cup E_{+}$, where ${E}_{+}$ and ${E}_{-}$ are finite or countable collections of semi-infinite edges $e$ parametrized by $(-\infty,0)$ and $(0,+\infty)$, respectively. The half-lines are connected at a unique vertex $\nu=0$. In this case the graph is sometimes called a star-shaped metric graph. Several arguments used in [44] and [38] in star-shaped graphs are extended in this work for looping-edge graphs. We recall that when the star-shaped metric graph is balanced ($|{E}_{+}|=|{E}_{-}|$), the extension theory implies that there is a $9|{E}_{+}|^{2}$-parameter family of skew-self-adjoint operators for $A_{0}$. Now, on looping-edge graph when the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we get that there must be $9(k+1)^{2}$ family of unitary groups, each one having its own representation.
@@
As far as we know, the study of the KdV equation on metric graphs remains relatively underdeveloped. Cavalcante [23] studied the local well-posedness of the Cauchy problem associated with (1.8) in Sobolev spaces $H^{s}(\mathcal{Y})$ with low regularity on a $\mathcal{Y}$-junction graph, $\mathcal{Y}=(-\infty,0)\cup(0,+\infty)\cup(0,+\infty)$. Subsequently, Angulo and Cavalcante [8] established a criterion for the linear instability of stationary solutions for the KdV on arbitrary star-shaped metric graphs. This criterion was applied to tail and bump profiles on balanced graphs with specific boundary conditions at the vertex $\nu=0$, ensuring they remain linearly unstable (see Theorem 6.2 in [8]). More recently, Mugnolo, Noja, and Seifert [39] investigated the existence of solitary waves on a star-shaped metric graph in terms of the coefficients of the KdV equation $(\alpha_{\mathbf{e}})_{e\in E}$, $(\beta_{\mathbf{e}})_{e\in E}$ on each edge, with coupling conditions at the vertex $\nu=0$, which may or may not involve continuity, and examined the speeds of the traveling waves. Lastly, Angulo and Cavalcante [10] developed a local well-posedness theory for the KdV on a balanced star graph, considering boundary conditions at the vertex $\nu=0$ of $\delta$- interaction type. They also extended the linear instability result from [8] to establish nonlinear instability.
Results of the dynamics of the KdV model on looping edge graphs or on $\mathcal{T}$-shaped graphs, such as local well-posedness of the Cauchy problem, existence and stability of solitary waves, etc., are not known in the literature to the best of the authors knowledge. We believe that the theory established in this manuscript can be considered as a first step towards a better understanding of this notable dispersive model on metric graphs. Organization of the article.
  1. 2.
    Preliminaries
    1. 2.1.
      Notation
    2. 2.2.
      Stone and Lummer-Phillips Theorems
    3. 2.3.
      Operators on Krein Spaces
  2. 3.
    The free Airy operator and deficiency indices
  3. 4.
    The Airy operator on a looping-edge graph
    1. 4.1.
      Boundary system for even amount of half-lines
      1. 4.1.1.
        Extensions generating unitary dynamics
      2. 4.1.2.
        Examples: $\delta$-type interactions on a looping-edge graph
    2. 4.2.
      Boundary system for arbitrary amount of half-lines
      1. 4.2.1
        Extensions generating contractile dynamics
      2. 4.2.2
        Examples: $\delta$-type interactions on a tadpole graph
    3. 4.3.
      Boundary system separating derivatives
      1. 4.3.1
        Extensions generating contractile dynamics via splitting derivatives
      2. 4.3.2
        Examples
  4. 5.
    The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped metric graphs.
    1. 5.1.
      Deficiency indices for the free Schrödinger operator
    2. 5.2.
      Boundary system
    3. 5.3.
      Extensions generating unitary dynamics
    4. 5.4.
      Extensions generating unitary dynamics parameterized by linear subspaces
    5. 5.5.
      Extensions to $\mathcal{T}$-shaped metric graphs

### v4

@@ Abstract
The aim of this work is to study the Airy and Schrödinger operators on looping-edge graphs, a class of metric graphs consisting of a circle and a finite number $N$ of infinite half-lines attached to a common vertex. For the Airy operator, we characterize all extensions generating unitary and contractive dynamics in terms of self-orthogonal subspaces and linear operators acting on indefinite inner product spaces (Krein spaces) associated to the boundary values at the vertex. Employing similar abstract techniques, we then describe a systematic way to produce self-adjoint extensions of the Schrödinger operator that are compatible with prescribed boundary relations on looping-edge and $\mathcal{T}$-shaped graphs.
@@ 1. Introduction
Nonlinear evolution models on metric graphs arise as models in wave propagation, for instance, in quasi-one-dimensional systems (e.g., meso- or nanoscale) that resemble a thin neighborhood of a graph. These models provide a convenient means to study various physical effects in the real world, both from theoretical and experimental perspectives. The flexibility in setting the geometry of the graph configuration allows for the creation of different dynamics. To mention a few examples of bifurcated systems: Josephson junction structures, networks of planar optical waveguides and fibers, branched structures associated with DNA, blood pressure waves in large arteries, nerve impulses in complex arrays of neurons, conducting polymers, and Bose-Einstein condensation (see [19, 21, 22, 26, 27, 29, 32, 38, 40, 43, 50] and references therein).
Recently, nonlinear models on graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries model, have been studied extensively (see [2, 13, 14, 10, 11, 15, 16, 17, 3, 32, 43] and references therein). From a mathematical viewpoint, an evolution model on a graph is equivalent to a system of PDEs defined on the edges (intervals), where the coupling is determined exclusively by the boundary conditions at the vertices (known as the “topology of the graph”), which governs the dynamics on the network via groups. This area of study has attracted significant attention, particularly in the context of soliton transport. Solitons and other nonlinear waves in branched systems provide valuable insights into the dynamics of these models.
@@
We recall that a metric graph $\mathcal{G}$ is a structure represented by a finite number of vertices $V=\{\nu_{i}\}$ and a set of adjacent edges at the vertices $E=\{e_{j}\}$ (for further details, see [19]). Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$. Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical “wires” or “networks” connecting them. The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$. This identification introduces the coordinate $x_{e}$ along the edge $e$.
An example of a metric graph, which will be the main focus of our study here, is a looping-edge graph. This is a graph composed of a circle and a finite number $N$ of infinite half-lines attached to a common vertex $\nu=L$. If the circle is identified with the interval [L,L][-L,L] and the half-lines with $L,\infty)$, we obtain a particular metric graph structure that we denote with $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$ and $V=\{L\}$ (see Figure 1). We use the subscripts 0 and $j$ to refer to the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$, $j=1,\dots,N$, respectively. The special case $N=1$ is known as a tadpole graph (see also Figure 1).
L-L
@@
A wave function $\mathbf{U}$ defined on $\mathcal{G}$ will be understood as an $(N+1)$-tuple of functions $\mathbf{U}=(\phi,(\psi_{j})_{j=1}^{N})=(\phi,\psi_{1},\dots,\psi_{N})$, where $\phi$ is defined on $e_{0}=[-L,L]$ and $\psi_{j}$ is defined on $e_{j}=[L,\infty)$ for $j=1,\dots,N$.
The primary aim of this manuscript is to develop a self-contained and systematic theory of linear dynamics for both the Airy and the Schrödinger operators on looping-edge graphs, intended to serve as a reference for future nonlinear applications. We emphasize that this abstract theory is presented as a standalone contribution: the dynamical part is established here independently, so that subsequent studies on existence, stability, controllability, or stabilization of nonlinear waves on these graphs can build upon it without having to reconstruct the linear theory from scratch.
For the Airy operator, to the best of our knowledge, no results on the characterization of linear dynamics on looping-edge graphs appear in the existing literature. We fill this gap by providing, in terms of Krein space techniques, a complete description of all boundary conditions at the vertex that make the Airy operator the generator of a unitary group or a contraction semigroup. For the Schrödinger operator, our approach offers an additional structural advantage compared to the classical methods: it allows one to prescribe a priori the type of vertex interactions desired (for instance, continuity of derivatives at the vertex) and then to systematically identify all self-adjoint extensions compatible with those conditions. In this way the theory does not merely produce a parameter family of extensions, but organizes them according to physically or geometrically motivated constraints.
@@
$i\mathbf{U}_{t}+\Delta\mathbf{U}+(p+1)|\mathbf{U}|^{2p}\mathbf{U}=\mathbf{0},\quad p>0,$ (1.1)
@@
$\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
@@
Since the deficiency indices of $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ are equal to $2+N$, the extension theory of Krein and von Neumann provides a $(2+N)^{2}$-parameter family of self-adjoint extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ of $\mathcal{H}_{0}$ on $L^{2}(\mathcal{G})$, where the action of $\Delta_{ext}$ is given by $\Delta_{ext}=-\Delta$. Each member of this family defines a unitary dynamic for the linear evolution problem
$\left\{\begin{array}[]{ll}i\mathbf{U}_{t}=-\Delta_{ext}\mathbf{U}\\ \mathbf{U}(0)=\mathbf{u}_{0}\in D(-\Delta_{ext}).\end{array}\right$. (1.4)
We point out that this approach, based on boundary systems [49], demonstrates that every extension operator $(-\Delta_{ext},D(-\Delta_{ext}))$ can be characterized by the boundary conditions at the vertices $\nu=\pm L$, treating the point L-L as a terminal vertex through a relatively simple matrix process (see Theorem 5.2 below). At this level, this part of the work can be seen as a complement to the works of Exner et al. [29, 30, 31], which, to our knowledge, has not formally appeared in the literature on $\mathcal{G}$. Furthermore, in general, the extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ do not necessarily satisfy continuity conditions at the boundary of [L,L][-L,L]. In the case of looping-edge graphs, the extensions must ensure that the wave functions are continuous at least on the loop (i.e., $\phi(L)=\phi(-L)$). For example, in the study of electron motion in thin metallic “wires,” it is required that the wave function remains continuous at the junction $\nu=L$. These boundary conditions have a clear physical interpretation, representing the conservation of probability flow at the junction (see [30]).
@@
$\begin{array}[]{ll}D_{Z,N}=\{\mathbf{U}\in H^{2}(\mathcal{G}):&\phi(L)=\phi(-L)=\psi_{1}(L)=\cdots=\psi_{N}(L)\;\text{and}\\ &\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{j=1}^{N}\psi_{j}^{\prime}(L+)+Z\psi_{1}(L)\},\end{array}$ (1.5)
where for any $n\geqslant 0$,
@@
The boundary conditions in (1.5) are called $\delta$-type if $Z\neq 0$, and Neumann–Kirchhoff type if $Z=0$. The parameter $Z$ is a coupling constant between the loop and the several half-lines. The choice of coupling at the vertex $\nu=L$ corresponds to a feasible quantum-wire experiment (see [29, 31] and references therein). It should be noted that, based on the domain $D_{Z,N}$, several recent studies have been conducted, particularly on the existence and stability of standing-wave solutions for the model in (1.1) (see [3, 4, 5, 7, 8, 18, 23, 24, 37, 46, 44] and references therein).
A distinctive feature of the approach developed here is that it allows one to impose vertex conditions a priori and then identify all extensions compatible with them, rather than parametrizing all extensions at once and then selecting the relevant subfamily. Concrete illustrations are provided in Examples 8 and 10 below. The existence and (in)stability of standing wave profiles for the NLS model (1.1) on looping-edge and $\mathcal{T}$-shaped graphs under dynamics induced by some of these extensions is an active research direction pursued by the authors. As a further illustration of this framework, in Section 5.7 we discuss the orbital instability of elementary (yet novel) standing waves for the cubic NLS on $\mathcal{G}$ under pure periodic plus Neumann type of interactions.
Another important goal of this work is to obtain a characterization of all boundary conditions under which the Airy-type evolution equation
$\partial_{t}u_{\mathbf{e}}(x,t)=\alpha_{\mathbf{e}}\partial_{x}^{3}u_{\mathbf{e}}(x,t)+\beta_{\mathbf{e}}\partial_{x}u_{\mathbf{e}}(x,t),\quad x\neq 0,\ t\in\mathbb{R},\ \mathbf{e}\in E,$ (1.6)
with real coefficients given by the sequences $(\alpha_{\mathbf{e}})_{e\in E}$, $(\beta_{\mathbf{e}})_{e\in E}$, generates either a contraction semigroup or a unitary group on a looping-edge graph (see Theorems 4.2, 4.3 and 4.6). Our approach is based on the abstract results in [49] applied to the case of the Airy operator
@@
when we consider the metric graph $\mathcal{G}$ represented by $E$ defined above and virtually treating the point L-L as a terminal vertex. In contrast with the symmetric operator case $\mathcal{H}_{0}$ the skew-symmetry of $A_{0}$ induce a slightly different boundary system and therefore different structure of the characterizations (compare Theorems 5.2 and 4.3 below) Therefore, notions of self-orthogonal subspaces and operators based on indefinite inner product spaces (Krein spaces) will be necessary. Thus, we obtain in particular a description of all (skew-)self-adjoint realizations of $A_{0}$, i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$. For instance, in Example 1 below we obtain $\delta$-interaction type boundary conditions on a looping-edge graph in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [10]) under conditions on $\alpha_{\mathbf{e}}$ and $\beta_{\mathbf{e}}$. We also establish conditions for obtaining extensions of $A_{0}$ generating contractive dynamics.
@@
We recall that the evolution linear model in (1.6) is associated to the following vectorial Korteweg–de Vries (KdV) equation
$\partial_{t}u_{\mathbf{e}}(x,t)=\alpha_{\mathbf{e}}\partial_{x}^{3}u_{\mathbf{e}}(x,t)+\beta_{\mathbf{e}}\partial_{x}u_{\mathbf{e}}(x,t)+2u_{\mathbf{e}}(x,t)\partial_{x}u_{\mathbf{e}}(x,t)$ (1.8)
@@
on the metric graph $\mathcal{G}$. Recently, Mugnolo, Noja, and Seifert [41] characterized all boundary conditions under which the Airy equation in (1.6) generates either a semigroup or a group on a metric graph that has a structure represented by the set $E\equiv E_{-}\cup E_{+}$, where ${E}_{+}$ and ${E}_{-}$ are finite or countable collections of semi-infinite edges $e$ parametrized by $(-\infty,0)$ and $(0,+\infty)$, respectively, connected at a unique vertex $\nu=0$. In this case the graph is sometimes called a star-shaped metric graph. Several arguments used in [49] and [41] in star-shaped graphs are extended in this work for looping-edge graphs. We recall that when the star-shaped metric graph is balanced ($|{E}_{+}|=|{E}_{-}|$), the extension theory implies that there is a $9|{E}_{+}|^{2}$-parameter family of skew-self-adjoint operators for $A_{0}$. On a looping-edge graph, when the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, there must be a $9(k+1)^{2}$-parameter family of unitary groups, each one having its own representation.
@@
As far as we know, the study of the KdV equation on metric graphs remains relatively underdeveloped. Cavalcante [25] studied the local well-posedness of the Cauchy problem associated with (1.8) in Sobolev spaces $H^{s}(\mathcal{Y})$ with low regularity on a $\mathcal{Y}$-junction graph, $\mathcal{Y}=(-\infty,0)\cup(0,+\infty)\cup(0,+\infty)$. Subsequently, Angulo and Cavalcante [10] established a criterion for the linear instability of stationary solutions for the KdV on arbitrary star-shaped metric graphs. This criterion was applied to tail and bump profiles on balanced graphs with specific boundary conditions at the vertex $\nu=0$, ensuring they remain linearly unstable (see Theorem 6.2 in [10]). More recently, Mugnolo, Noja, and Seifert [42] investigated the existence of solitary waves on a star-shaped metric graph in terms of the coefficients of the KdV equation $(\alpha_{\mathbf{e}})_{e\in E}$, $(\beta_{\mathbf{e}})_{e\in E}$ on each edge, with coupling conditions at the vertex $\nu=0$, which may or may not involve continuity, and examined the speeds of the traveling waves. Lastly, Angulo and Cavalcante [12] developed a local well-posedness theory for the KdV on a balanced star graph, considering boundary conditions at the vertex $\nu=0$ of $\delta$-interaction type, and extended the linear instability result from [10] to establish nonlinear instability.
Results on the dynamics of the KdV model on looping-edge graphs, such as local well-posedness of the Cauchy problem and existence and stability of solitary waves, are not known in the literature to the best of the authors’ knowledge. The theory developed in this manuscript is intended to fill that gap at the linear level and to lay the groundwork for future nonlinear investigations. Several applications are already emerging from the framework presented here. On the one hand, the boundary system based on separation of first derivatives, developed in Section 4.3, has already proven effective in the analysis of dispersive equations on metric graphs: the technique was used in [42] to construct explicit solitary-wave solutions for the KdV equation on star graphs, where the first-derivative coupling at the vertex plays a distinguished role. Moreover, the domains generated via separation of derivatives in Example 7 directly recover the looping-edge analogue to the vertex conditions employed in [6, 48], works on the stabilization and controllability of the KdV on networks with bounded and unbounded domains. In general these abstract theory can be used to yield the class of boundary conditions needed for controllability problems on metric graphs, in fact the class of extensions $A_{Y,R_{1}}$ constitute the functional-analytic backbone of an ongoing program on exponential stabilization and controllability of the KdV equation on $\mathcal{G}$, currently in an advanced stage of research. As a concrete illustration of the use of contractive dynamics, in Section 4.4 we show that adding a damping term $\gamma\mathbf{U}$ to the Airy evolution under the boundary conditions of Example 6 yields exponential decay of the $L^{2}(\mathcal{G})$-energy at rate $\gamma$, with the dissipation arising entirely from the boundary structure rather than from the dispersive coefficients.
Organization of the article. Section 2 collects the necessary preliminaries, including notation, the Stone and Lummer–Phillips theorems, and the relevant theory of operators on Krein spaces. Section 3 analyzes the free Airy operator and computes its deficiency indices on the looping-edge graph. Section 4 is the core of the paper and is devoted to the Airy operator on looping-edge graphs. It is divided into three parts: Section 4.1 treats the case of an even number of half-lines and characterizes extensions generating unitary dynamics; Section 4.2 handles an arbitrary number of half-lines and characterizes extensions generating contractive dynamics; Section 4.3 develops an alternative boundary system based on the separation of derivatives, providing a further class of extensions with contractive dynamics. Finally, Section 5 applies the same abstract framework to the Schrödinger operator on looping-edge graphs, characterizing its self-adjoint extensions and the associated unitary dynamics.